A research project you are working on requires that an electron have a wave function represented by

𝜓(x) = Aebx for x ≥ 0Ae+bx for x < 0

The experiment requires that the electron have a 98.0% chance of being within a distance of a = 14.5 cm

 from x = 0,

 on either side of this center point. From this information, you need to determine numerical values for b (in m−1) and the normalization constant A (in m−1/2) for the wave function.


1
Expert's answer
2021-05-09T13:03:38-0400

The probability of finding electron from x = 0 to x = 14.5 cm "= 14.5 \times 10^{-2} \;m" is P = 98 %

The probability of finding electron:

"P = \int_{x=0}^{0.145} \psi^*(x) \psi(x)dx = 98 \; \% \\

P = \int_{x=0}^{0.145}(\sqrt{b}e^{-bx})^* \sqrt{b}e^{-bx}dx = 98 \; \% \\

\int_{x=0}^{0.145} be^{-2bx}dx = 98 \; \% \\

[\frac{b \times e^{-2bx}}{-2b}]^0.145_0 = 98 \; \% \\

[ e^{-2bx}]^0.145_0 = \frac{98}{100} \times (-2) \\

[ e^{-2b \times 0.145 – e^{-0}}] = \frac{-98}{50} \\

e^{-0.29b} -1 = \frac{-98}{50} \\

e^{-0.29b} = 1 -\frac{98}{50} \\

e^{-0.29b} = \frac{50 -98}{50} \\

e^{-0.29b} = \frac{-48}{50} \\

e^{0.29b} = \frac{-50}{48} = -1.0416 = e^{0.041} \\

b = \frac{0.041}{0.29} = 0.141 \;m^{-1} \\

A = \sqrt{b} = \sqrt{0.141} = 0.376"